What are the equations for motion regarding distance, velocity, acceleration. Also, please mention that the equations would be if the object was in free fall
Answer #1

From equation of motion

Y = v_{0y} t +\frac{1}{2} a_y t^2

distane Y = \frac{1}{2} gt^2

from equation of motion

v = u + at

velocity v = g t

incase of acceleration

a = g